40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log, Page 155 of 328  Not logged in ELOG logo
ID Date Author Type Category Subject
  8701   Thu Jun 13 16:01:48 2013 nicolasUpdateGreen LockingY arm locked with green but bad mode matching

Quote:

After restoring alignment I could see again strong 00 flashes (about 250-300 counts on ALS-TRY). So I locked the arm with IR and after enabling the PDH servo for the green locking, I also locked the green on the Y arm in 00 mode. Then I moved the two mode matching lenses to maximize the power into the 00 mode, but I didn't reach more than 30-35 counts.

Green power injected into the Y arm                    0.680mW

Green power reflected back                                  0.090mW

Green power transmitted on the PSL                  few uW

I would expect more power on the PSL table (maybe 10x more).

Is this reflection measured with the cavity locked or unlocked?

So what's the actual designed reflectivity of the ETM for green? No one seems to be able to give me a straight answer about this.

Looking at the reflected beam when the beam is misaligned makes it look like it's << 0.9. Is that expected given the coating spec?

You say the cavity scan goes as high as 300cts but you can only lock to 30cts, are you locked on the sideband?

  8700   Thu Jun 13 15:04:16 2013 JenneUpdateLSCNew modeled sensing matrix

Using all of the latest parameters that I can find, I have re-modeled the 40m sensing matrix.  Also, I have it output the data in a format that can be used by the same plotting function as the measured sensing matrix, so they are nice and easy to compare.

The newly modeled 40m sensing matrix:

SensMatModel_13June2013.png

To compare, here is the measured sensing matrix from elog 8644:

SensMatMeas_23May2013.png

Notice that (a) the units are different, so don't focus too much on the amplitudes of the lines, and (b) all of the measured and modeled matrix elements are pretty similar, except for the REFL11.  REFL11 (top right in model plot, top center in measured plot) looks like it's flipped, as well as rotated.  The new model doesn't match up too well with the Kiwamu/Koji models (which matched eachother okay), but I like that the new model matches the measurements fairly well.  The Koji sensing matrix: on the 40m wiki

EDIT: I have replaced the modelled plot with a new version.  The data and numbers are the same, but I have switched the labels on the individual radar plots, and forced them to be in the same order as they are in the measured plot.

  8699   Thu Jun 13 10:57:36 2013 AnnalisaUpdateGreen LockingY arm locked with green but bad mode matching
> > Hmmm.  You seem to be saying that more light is reflected than is injected. Is this a units problem? Or was some IR on the power meter during the 'reflected' measurement? 
> > We should look at it with fresh eyes in the morning. 
> 
> Also, if you have been measuring the power of green refl at the rejection port of the green faraday, the polarization of the light entering the green faraday should be checked once again to make sure that you are measuring 
> only the reflected power from the arm cavity.

Sorry Sorrry Sorry!!
It was 0.090 mW, I just forgot a zero!!!
Sorry!
  8698   Thu Jun 13 00:35:23 2013 ManasaUpdateGreen LockingY arm locked with green but bad mode matching
> Hmmm.  You seem to be saying that more light is reflected than is injected. Is this a units problem? Or was some IR on the power meter during the 'reflected' measurement? 
> We should look at it with fresh eyes in the morning. 

Also, if you have been measuring the power of green refl at the rejection port of the green faraday, the polarization of the light entering the green faraday should be checked once again to make sure that you are measuring 
only the reflected power from the arm cavity.
  8697   Wed Jun 12 23:06:33 2013 JenneUpdateGreen LockingY arm locked with green but bad mode matching
Hmmm.  You seem to be saying that more light is reflected than is injected. Is this a units problem? Or was some IR on the power meter during the 'reflected' measurement? 
We should look at it with fresh eyes in the morning. 
  8696   Wed Jun 12 22:48:10 2013 AnnalisaUpdateGreen LockingY arm locked with green but bad mode matching

After restoring alignment I could see again strong 00 flashes (about 250-300 counts on ALS-TRY). So I locked the arm with IR and after enabling the PDH servo for the green locking, I also locked the green on the Y arm in 00 mode. Then I moved the two mode matching lenses to maximize the power into the 00 mode, but I didn't reach more than 30-35 counts.

Green power injected into the Y arm                    0.680mW

Green power reflected back                                  0.090mW

Green power transmitted on the PSL                  few uW

I would expect more power on the PSL table (maybe 10x more).

  8695   Wed Jun 12 01:56:58 2013 AnnalisaUpdateGreen LockingY arm locked with green but bad mode matching

 

 For the mode matching calculation I was using the ETMY focal length that I found on Kiwamu's plot on the wiki page. 

Taking into account also the substrate, the focal length turns out to be

fl = ((n-1)*(1/R1 - 1/R2 + (n-1)d/(nR1R2)))^(-1) = -125.81 m

with n = 1.46071 (refraction index of fused silica at 532nm)

R1 = 5625 m (radius of curvature of the first surface)

R2 = 57.37 m (radius of curvature of the second surface)

d = 25mm (thickness)

 

The value of the focal length is sligthly different from the one I was using before in the calculation, but maybe it is enough to change the coupling.

The mode matching solution I found is very sensitive to the lenses position. 

The beam waist position can vary up to 20m varying by 1cm the first lens position, while it is slightly less sensitive to the second lens displacement.

As shown in the picture, along the green beam path there is also a 1m focal length lens. It's position is fixed, because it is along the IR transmetted beam path also. I tried to get a better solution without it, but I found that the waist position was still strongly dependent on one of the two lenses position, so it would not solve the problem to remove this lens.

I think that the main issue of this mode matching is related to the "space contraints", because the two lenses' positions can vary in a very small space, even though the green beam path on the table is quite long.

Eventually, I put the MM lenses found from this last simulation on the table, and it seems to work, since I've seen very strong 00 flashes. Unfortunately, while trying to maximize the alignment I broke it  and I have to do it again, but I feel confident! 

Attachment 1: waistVaryingLensPos.pdf
waistVaryingLensPos.pdf
Attachment 2: green.JPG
green.JPG
Attachment 3: ModeMatchingY.pdf
ModeMatchingY.pdf
  8694   Tue Jun 11 22:16:56 2013 ManasaSummaryGreen LockingALS for X arm

I discussed with Yuta about the ALS servo and phase tracker and found that there was a lot of information lying around from last year but there aren't any clear elogs on how to enable ALS and obtain IR resonance.

 

Guide to enabling the ALS servo and find IR resonance:

The steps will explain in detail how to ressurrect the ALS servo for green X-arm and find IR resonance using ALS. The medm screens are very confusing right now.

 

(i) Finding the beat note

1. Get the IR to flash in TEM00 for the arm and lock it by enabling LSC (Locking the arm to IR keeps the arm cavity mirrors stable so that you can scan the temperature of the X-end laser to find the beat note).

2. Steer the X-green into the arm cavity such that the arm cavity locks in TEM00 for green as well. At this point you should also have the X-green reaching the PSL table.

3. Align the PSL doubled green (PSL-green) and the X-green in near-field (at the camera) and far-field (letting the beams to propagate beyond the Green-TRX PD).

4. Check cabling of the RF beat PD.

5. Change the X-laser temperature by sweeping the offset (C1: ALS-SLOW_SERVO2_OFFSET) in steps of 10.

6. Find the beat note and tune the alignment at the beat PD to maximize the beatnote amplitude. Disable LSC for X arm.

 

(ii) The GREEN HORNET explained

'Input signal conditioning' block takes I and Q signals after the delay frequency discriminator (DFD) in the beat box and these signals pass through C1ALS_BEATX_FINE filter banks. The output signal then enters the phase rotation matrix of the phase tracker. The phase tracker gives 'PHASE_OUT' which is the error signal that is fed to the ETM servo filter module (DOF filters)  through the 'Input matrix' in the medm. 

An offset can also be fed to the phase tracker which will scan the beat frequency (used to find IR resonance).

 

(iii) Scripts

1. easyALS.py - This runs from 'ON plus' or 'ON minus' buttons in the C1ALS_COMPACT. 

The script clears history of 'fine_phase' filter module and increases gain of the servo in steps ('ON plus' for positive gain and 'ON minus' for negative gain).

2. findIRresonance.py - This runs from 'IRres' button in the C1ALS_COMPACT.

It adds offset to the phase tracker in steps which scans the beat frequency to find IR resonance.

P.S. Check the scripts before enabling the servo so that the right filter modules are being turned ON. Using the wrong set of filter modules can kick the ETM.

____________________________________________________________________________________________________________________________________________________

X arm ALS progress:

I found the beat note and got ALS to work reasonably for the Xarm without kicking the ETM. I did this by manually toggling buttons and changing gains. The scripts need editing.

To do:

Modify the scripts to work as we want them to.

The ALS medm is SSSOOOO confusing. It definitely needs to be fixed (remove all unwanted parts of the screen that existed 'pre-phase tracker').

Find IR resonance.

 
  8693   Tue Jun 11 10:00:54 2013 nicolasUpdateGreen LockingY arm locked with green but bad mode matching

Quote:

 

Still no good locking!

After making the reflected beam size closer to the injected one, I maximized alignment. I locked again in 00 mode, but I couldn't maximize the power. 

I just realized that maybe I'm not using the correct radius of curvature for the ETMY in the simulation. Tomorrow I will start checking from that.

 Also make sure you are taking into account the substrate of the ETM.

  8692   Mon Jun 10 21:39:26 2013 AnnalisaUpdateGreen LockingY arm locked with green but bad mode matching

 

Still no good locking!

After making the reflected beam size closer to the injected one, I maximized alignment. I locked again in 00 mode, but I couldn't maximize the power. 

I just realized that maybe I'm not using the correct radius of curvature for the ETMY in the simulation. Tomorrow I will start checking from that.

  8691   Sat Jun 8 00:24:11 2013 AnnalisaUpdateGreen LockingY arm locked with green but bad mode matching

[Annalisa, Nic]

After connecting the PD with the reflection from the arm  to the PDH box, theY  arm has been locked on the 01 mode.  Maximizing the alignment, we obtained a 00 mode locking, but we couldn't maximize the power.

The size of the reflected beam was different with respect to the size of the incoming beam, so probably a bad mode matching was one of the issues.

Moreover, the reflected beam is very low power. We need to figure out why it is so (bad alignment? related to mode matching?)

 

After measuring better all the distances, I did a new mode matching calculation. I put the lenses after measuring the beam waist, so the size of the beam on the lenses was the same as expected from the calculation. Nevertheless, the beam size on the beam splitter looks bigger than expected, and also in this case green flashes into the cavity at some HOM (again 01).

I also tried to lock again the cavity and maximize the alignment, but I didn't get any improvement with respect to the previous mode matching.

 

  8690   Fri Jun 7 23:44:54 2013 KojiUpdateGeneralProjector - lightbulb replaced

http://nodus.ligo.caltech.edu:8080/40m/7885

  8689   Fri Jun 7 21:33:39 2013 ManasaUpdatePSLPSL green to Y- BEAT PD blocked

To allow Annalisa to work on the Y-green alignment as I work with the X-green, the part of the PSL green beam that goes to the Y-green beat PD has been blocked with an iris.

  8688   Fri Jun 7 17:12:46 2013 gautamUpdateGeneralProjector - lightbulb replaced

Quote:

Quote:

Update: We don't have our BIG screen 

There was no light from the projector when I came in this morning. I suspected it might have to do with the lifetime of the bulb. But turning the projector OFF and ON got the projector working....but only for about 10-15 seconds. The display would go OFF after that. I will wait for some additional help to dismount it and check what the problem really is.

 

 

  • Shipping Estimate Friday June 7, 2013 - Monday June 10, 2013
    Delivery Estimate: Wednesday June 12, 2013 - Monday June 17, 2013 by 8:00pm
  • Lampedia Projector Lamp for VIEWSONIC PRO8200 / PRO8300
    Sold by: Lampedia

 [manasa, gautam]

-the replacement lamp arrived a while back.

-the old lamp has been switched out, it had 3392 lamp hours on it.

-new lamp installed, projector mounted back up, and lamp hours reset to zero. there is a lingering odour of something burning, not sure what it is or if it is in any way connected to the new lamp. old lamp disposed in the hazardous waste bin. the big screen is back online.

  8687   Thu Jun 6 16:33:19 2013 steveUpdatesafetyupdated laser inventory

 Updated laser inventory and operator list. They are posted in the 40m wiki and entry doors of the 40m IFO room.

Let me know if this list needs correction.

Lightwave  M126N-1064-700 NPRO sn 337 in the PSL enclosure got connected to local emergency shut off switch. This is a LIGO operational safety requirement.

 

Attachment 1: ELSOconnected.jpg
ELSOconnected.jpg
  8686   Thu Jun 6 15:46:10 2013 Max HortonSummaryGeneral Smart UPS 2200 Batteries Replaced

Replaced the batteries successfully in the control room.  We just had to switch the clips from the old batteries to the new one, which we didn't know was possible until now.

  8685   Thu Jun 6 09:58:13 2013 SteveUpdateendtable upgradeenclosure to chamber connection

 

 Thin wall connector from McMCarr#55275K25 was tested in 150 mW, 1 mm beam size of 1064 nm  overnight. It did not show any degradation.

Super-Compressible Duct for Air

Hose is made from 0.005" thick, double-ply metalized polyester with a fabric-enclosed steel wire support. 

Atm2, Enclosure viewport adaptor is shown in place of the viewport.

Soft gaskit -   durumeter hardness 10A - McMCarr#9010K51 was added on the 10" od sufaces of conflat and viewport adaptor to insure being air tight.

The duct connector clamped with soft braided elastic " Velstrech" brand loop.

 

Attachment 1: thinwallconnector.jpg
thinwallconnector.jpg
Attachment 2: 06061301.PDF
06061301.PDF 06061301.PDF
  8684   Wed Jun 5 20:45:19 2013 gautamUpdateGeneralITMx Oplev Fixed

Quote:

 (Manasa)

The ITMx Oplev was misaligned. Switched the ITMx Oplev back on and fixed the alignment.

EDIT, JCD:  This is totally my fault, sorry.  I turned it off the other day when I was working on the POP layout, and forgot to turn the laser back on.  Also, I moved the fork on the lens directly in front of the laser (in order to accommodate one of the G&H mirrors), and I nudged that lens a bit, in both X and Y directions (although very minimally along the beam path).  Anyhow, bad Jenne for forgetting to elog this part of my work.

 [Jenne, Gautam]

It turned out that the earlier fix was not really a fix, because there was some confusion as to which of the two lenses Jenne moved while working, and while Manasa and I were re-aligning the beam, we may have moved the other lens.

Subsequently, when we checked the quadrant sum, it was low (in the region of 20), even though OPLEV_PERROR and OPLEV_YERROR were reasonably low. We called up a 30 day trend of the quadrant sum and found that it was typically closer to 4000. This warranted a visit to the table once again. Before going to the table, we did a preliminary check from the control room so as to make sure that the beam on the QPD was indeed the right one by exciting ITMx in pitch (we tried offsets of 500 and -500 counts, and the spot responded as it should). ITMx oplev servo was then switched off.

At the table, we traced the beam path from the laser and found, first, that the iris (I have marked it in one of the photos attached) was practically shut. Having rectified this, we found that the beam was getting clipped on the first steering mirror after the laser (also marked in the same photo, and a second photo showing the clipping is attached). The beam isn't very well centred on the first lens after the laser, which was the one disturbed in the first place. Nevertheless, the path of the entering beam seems alright. The proposed fix, then, is as follows;

  1. Use the existing iris and a second one to mark the path of the entering beam.
  2. Make the necessary adjustments (i.e. move the lens immediately after the laser and the first steering mirror after the laser) such that the direction of the entering beam is unchanged, and the clipping and centering problems are resolved. 

Back in the control room we noticed that the quadrant sum had gone up to ~3500 after opening out the iris. The OPLEV_PERROR and OPLEV_YERROR counts however were rather high (~200 counts in pitch and ~100 counts in yaw). Jenne went back to the table and fixed the alignment such that these counts were sub-10, and the quadrant sum went up to ~3800, close to the trend value. 

At the time of writing, the beam is still not centred on the lens immediately after the laser and is still getting clipped at the first steering mirror. Oplev servo back on.

Photos

 An overview

ITMx_overview.JPG

 

The clipping

ITMx_clipping_.JPG

  8683   Wed Jun 5 17:37:10 2013 ranaUpdateGeneralfast 2004 qpd

Quote:

Thank you Ben Abbott forwarding this information:

 QPD Amplifier D990272 https://dcc.ligo.org/cgi-bin/private/DocDB/ShowDocument?.submit=Number&docid=D990272&version= at the X-end.  It plugs into a Generic QPD Interface, D990692, https://dcc.ligo.org/cgi-bin/private/DocDB/ShowDocument?.submit=Number&docid=D990692&version= according to my drawings, that should be in 1x4-2-2A.

 Wrong: this is not an interface.

  8682   Wed Jun 5 15:55:16 2013 SteveUpdateGeneralProjector - lightbulb ordered

Quote:

Update: We don't have our BIG screen 

There was no light from the projector when I came in this morning. I suspected it might have to do with the lifetime of the bulb. But turning the projector OFF and ON got the projector working....but only for about 10-15 seconds. The display would go OFF after that. I will wait for some additional help to dismount it and check what the problem really is.

 

 

  • Shipping Estimate Friday June 7, 2013 - Monday June 10, 2013
    Delivery Estimate: Wednesday June 12, 2013 - Monday June 17, 2013 by 8:00pm
  • Lampedia Projector Lamp for VIEWSONIC PRO8200 / PRO8300
    Sold by: Lampedia
  8681   Wed Jun 5 15:48:02 2013 SteveUpdateGeneralfast 2004 qpd

Thank you Ben Abbott forwarding this information:

 QPD Amplifier D990272 https://dcc.ligo.org/cgi-bin/private/DocDB/ShowDocument?.submit=Number&docid=D990272&version= at the X-end.  It plugs into a Generic QPD Interface, D990692, https://dcc.ligo.org/cgi-bin/private/DocDB/ShowDocument?.submit=Number&docid=D990692&version= according to my drawings, that should be in 1x4-2-2A.

Attachment 1: D990272qpd.jpg
D990272qpd.jpg
Attachment 2: interfaceD990692.jpg
interfaceD990692.jpg
  8680   Wed Jun 5 15:03:42 2013 AnnalisaUpdateLSCC1:ALS-TRY_OUT filter and green progress

Quote:

The rejected beam from this Faraday comes out at a tiny, tiny angle and so its tough to pick it off without clipping the main beam.

Some care must be taken in setting this up - Steve may have some good ideas on what kind of mount can be placed so close to the beam.

Why did we ever order this terrible Faraday? Let's never get a Faraday with a tiny angle between the beams again.

The rejected beam from the Faraday is steered with a mirror into the PDA32A PD  and a 75mm fl lens is used to focus the beam into it.

The main beam is a few millimeters away from the mirror mount (maybe 2mm), and I think it should be fine as long as the main beam is not supposed to move.

 

Attachment 1: faraday_rejected_beam.JPG
faraday_rejected_beam.JPG
  8679   Wed Jun 5 14:43:42 2013 AnnalisaUpdateLSCALS-TRY_OUT DQ channels

Quote:

 After working some more on the EY table, we are getting some TEM00 flashes for the Y arm green. We have had to raise the height of one of the MM lenses to prevent clipping.

We used a function generator to apply a ~300 mV 10 Hz triangle wave to scan the laser frequency while aligning.

We tried to use the C1:ALS-TRY_OUT channel to help us in our alignment but there are a couple problems:

1) It seems that there is an uncompensated whitening filter before the ADC - Annalisa is making a compensation filter now.

2) The data delay is too much to use this for fast alignment. We might need to get a coax cable down there or mount a wired ethernet computer on the wall.

3) We need to make DQ channels for the TRY and TRX OUT. We need long term data of these, not just test points.

 [Jenne, Annalisa]

DQ channels have been created in the C1ALS model for TRX and TRY. They are called TRX_OUT and TRY_OUT and the sampling rate is 2048 Hz.

  8678   Wed Jun 5 14:39:41 2013 Max HortonUpdate Importing New Code

Most dependencies are met.  The next issue is that matplotlib.basemap is not installed, because it is not available for our version of python.  We need to update python on megatron to fix this.

  8677   Wed Jun 5 14:01:37 2013 ranaUpdateLSCC1:ALS-TRY_OUT filter and green progress

The rejected beam from this Faraday comes out at a tiny, tiny angle and so its tough to pick it off without clipping the main beam.

Some care must be taken in setting this up - Steve may have some good ideas on what kind of mount can be placed so close to the beam.

Why did we ever order this terrible Faraday? Let's never get a Faraday with a tiny angle between the beams again.

  8676   Wed Jun 5 10:46:43 2013 GautamUpdateGeneralCable re-routing at 1Y4

 There were 4 cables running over the front side of rack 1Y4 such that the front door could not be closed. I re-routed them (one at a time) through the opening on the top of the rack. The concerned channels were

  • Green refl mon
  • Err mon
  • Pzt out (temp) (has been marked "Door Damaged BNC")
  • Laser temp ctrl

Before and after pics attached.

Before

.before.JPG

After

after.JPG

 

 

 

  8675   Wed Jun 5 10:22:12 2013 ManasaUpdateGeneralProjector - viewsonic

Update: We don't have our BIG screen 

There was no light from the projector when I came in this morning. I suspected it might have to do with the lifetime of the bulb. But turning the projector OFF and ON got the projector working....but only for about 10-15 seconds. The display would go OFF after that. I will wait for some additional help to dismount it and check what the problem really is.

 

  8674   Tue Jun 4 21:50:23 2013 AnnalisaUpdateLSCC1:ALS-TRY_OUT filter and green progress

 

 [Annalisa, Gautam]

 The green beam alignment has been improved, so we see much more 00 bright flashing. We checked the polarization and the Ygreen shutter is back in place.

A mirror is already in place to steer the rejected beam from the green Faraday into a PD, tomorrow morning we'll put a lens and the PD to take the signal for PDH locking.

 

  8673   Tue Jun 4 20:19:07 2013 ranaConfigurationIOOChanged threshold for FSS SLOW loop

The FSS SLOW actuator often runs off away from zero and into a region where the mode hopping is bad and makes a lot of frequency noise (e.g. that 8 hour period a few weeks ago when Jamie couldn't lock the MC).

It should not have this behavior. The SLOW loop should only be running when the MC is locked.

Today I found that the threshold was set back to 0.2 V (which is approximately the correct value for the RefCav locking). Its being compared to the MC TRANS, so the correct value should be ~1/2 of the maximum MC TRANS.

To find this out, I read this piece of text:

    # Make sure the loop is supposed to be active
    if (get_value("C1:IOO-MC_TRANS_SUM") < get_value("C1:PSL-FSS_LOCKEDLEVEL")) {
    print("Reference Cavity not locked -- control loop disabled.\n");
    next;
    }

from scripts/PSL/FSS/FSSSlowServo. I set the threshold using the commmand:

caput -c -t C1:PSL-FSS_LOCKEDLEVEL 12000

Then I restarted the code on op340m, by typing:

> nohup FSSSlowServo

and then closing the terminal. Seems to be behaving correctly now. Previously, the value was so low that the SLOW loop was never turning itself off.

  8672   Tue Jun 4 17:28:09 2013 AnnalisaUpdateLSCC1:ALS-TRY_OUT filter and green progress

 [Annalisa, Gautam, Rana]

I made the anti-whitening filter for the C1:ALS-TRY_OUT channel.

zpk [[150],[15],1] Hz

Now we can look at the picks of this signal to align the green into the cavity.

We already had some 00 flash, but a better alignment has to be done.

TO DO:

- put the shutter along the beam path

- check the polarization (we have a new PBS for visible)

 

Attachment 1: green.JPG
green.JPG
Attachment 2: QUAD1_1054358327.mp4
  8671   Tue Jun 4 16:04:34 2013 AnnalisaUpdateLSCALS-TRY_OUT needs anti-whitening

Quote:

 After working some more on the EY table, we are getting some TEM00 flashes for the Y arm green. We have had to raise the height of one of the MM lenses to prevent clipping.

We used a function generator to apply a ~300 mV 10 Hz triangle wave to scan the laser frequency while aligning.

We tried to use the C1:ALS-TRY_OUT channel to help us in our alignment but there are a couple problems:

1) It seems that there is an uncompensated whitening filter before the ADC - Annalisa is making a compensation filter now.

2) The data delay is too much to use this for fast alignment. We might need to get a coax cable down there or mount a wired ethernet computer on the wall.

3) We need to make DQ channels for the TRY and TRX OUT. We need long term data of these, not just test points.

 I made the anti-whitening filter for the C1:ALS-TRY_OUT channel. But then I forgot to make an ELOG because I am bad.

  8670   Tue Jun 4 10:47:57 2013 SteveUpdatesafetysafety training

Our early bird surf student Gautem has received 40m specific basic safety training today.

  8669   Tue Jun 4 10:44:13 2013 SteveUpdateendtable upgradePI pzt holders are ready

The PI pzt holders are back from the shop. They are numbered 1, 2 & 3 and machined to match.

Tapered black delrin opener is to gauge the gap if it is too small to fit pzt. This is to prevent holder to be opened too much.

Attachment 1: PIpztholders.jpg
PIpztholders.jpg
  8668   Tue Jun 4 10:33:09 2013 ranaUpdateSUSETMY oplev path redone

 We also redid the oplev path.

Someone had used the illegal tactic of using aluminum dogs to hold down mounts which already have screw holes in the bases. This is too flimsy to use.

We then adjusted the position of the second lens to get a ~1 mm spot on the QPD. There is a lot of stray red light from extra reflections, but so be it.

Its possible that we are now clipping the IPANG path, but we will need to lock the Y arm and verify the input pointing to be sure. At this point we are ready to measure the ETMY OL loop gain and tune, etc.

  8667   Mon Jun 3 22:36:07 2013 ranaUpdateLSCALS-TRY_OUT needs anti-whitening

 After working some more on the EY table, we are getting some TEM00 flashes for the Y arm green. We have had to raise the height of one of the MM lenses to prevent clipping.

We used a function generator to apply a ~300 mV 10 Hz triangle wave to scan the laser frequency while aligning.

We tried to use the C1:ALS-TRY_OUT channel to help us in our alignment but there are a couple problems:

1) It seems that there is an uncompensated whitening filter before the ADC - Annalisa is making a compensation filter now.

2) The data delay is too much to use this for fast alignment. We might need to get a coax cable down there or mount a wired ethernet computer on the wall.

3) We need to make DQ channels for the TRY and TRX OUT. We need long term data of these, not just test points.

Attachment 1: try.pdf
try.pdf
  8666   Mon Jun 3 21:09:55 2013 ManasaUpdateGreen LockingX green beatnote through beatbox

I connected the X green beat PD output back to the beatbox, did the usual PSL alignment for green and searched for the beat note from the RFmon of the beatbox.

Yuta had used a power splitter which took Xbeat-RFmon and Ybeat-RFmon and used the SUM port to monitor the beat signals. I have removed this splitter and just used the X beat RFmon.

I found the beat note with:

Beat@58.7MHz : Amplitude -30dBm
C1:ALS-TRX_OUT16 = 3000 counts
C1:PSL-FSS_SLOWDC = 0.2250
PSL temperature = 31.52 degC
X- green temperature = 39.34 degC (OFFSET = 5140)

Next
Beatbox calibration

IMG_0608.JPG

  8665   Mon Jun 3 18:49:44 2013 GautamUpdateGeneralITMx Oplev Fixed

 (Manasa)

The ITMx Oplev was misaligned. Switched the ITMx Oplev back on and fixed the alignment.

EDIT, JCD:  This is totally my fault, sorry.  I turned it off the other day when I was working on the POP layout, and forgot to turn the laser back on.  Also, I moved the fork on the lens directly in front of the laser (in order to accommodate one of the G&H mirrors), and I nudged that lens a bit, in both X and Y directions (although very minimally along the beam path).  Anyhow, bad Jenne for forgetting to elog this part of my work.

  8664   Mon Jun 3 09:46:40 2013 SteveUpdateVACRGA scan at day 100

 Vacuum normal condition at pumdown day 100

 

Attachment 1: d100Mpd75.png
d100Mpd75.png
Attachment 2: d100Mpd75p.png
d100Mpd75p.png
Attachment 3: badguys.png
badguys.png
  8663   Sat Jun 1 14:14:56 2013 JenneUpdateASCNew POP path

I have a lens solution for the new POP QPD, plotted below.  To get the beam size, I started with the waist at the ITM, so the out of vacuum table starts around 6 meters on this plot.  Also, "PD" is the QPD, but the position marked on the plot is the maximum distance from the 2nd lens.  In reality, I will place it a few cm after the lens.  Once I've got that laid out, I'll move the 110PD (and its lens) and the camera around so that they are in good spots relative to the beam size.

 POP_QPD_lensSolution.png

Here is a photo of the way I left the table last Thursday.  The notations in orange indicate what I need to do to make the actual table match my lens solution.

 POX_table_30May2013progress.pdf

 

  8662   Fri May 31 11:15:01 2013 AnnalisaUpdateendtable upgradeETMY - Mode Matching for green - new calculation

 Since the beam waist after the Faraday had changed since the last time I measured it (maybe alignment changed a bit), I made a new mode matching calculation for green. I attached the results.

I'm going to align the beam into the Yarm.

RXA: JPG images deleted - replace with PDF please.

  8661   Fri May 31 10:25:17 2013 steveUpdatesafetyfirst aid kits refilled

Quote:

 

                                                 Recommended correction list:

1,  refill- upgrade first aid boxes

2,  maintain 18" ceiling to bookshelf clearance so the ceiling fire sprinklers are not blocked: room 101

3,  label chilled water supply & return valves in IFO room

4,  calibrate bake room hoods annually

5,  update safety sign at fenced storage

 

              40m still to do list:

1,   clean and measure all safety glasses

2,   annual crane inspection is scheduled for 8am March 19, 1013

3,   make PSL encloser shelf earthquake proof

 

Do you see something that is not safe? Add it to this list please.

 

 

 

Restocked First Aid Kits Location:

Main entrance, room 100

Drill press - above N2 cylinders, room 103

Control room, next to fire extinguisher, room  102

Vertex-north wall, IFO room 104

ETMY - right on ends light switches, IFO room 104_ east end

ETMX - on vertical I-beam of crane, IFO room 104_south end

Behind 1X3 Rack, on south wall - under instrument breakers panel PC-1, IFO room 104

 

Last thing remaining to be fixed from 2013 Safety Audit:

replace book shelf with 83" height

 

 

Attachment 1: refilledFAidK.jpg
refilledFAidK.jpg
  8660   Thu May 30 20:45:53 2013 AnnalisaUpdateendtable upgradeETMY - oplev

Quote:

Quote:

 Temporary oplev in place. The spot on the qpd is still big. My two lens solution did not work.

I will finalize optical component position of the oplev after the the arm transmitted and green beam optics in place. They have priority.

 

 Oplev spot size on qpd ~ 1 mm 

PS: I realized it later that the returning beam is going through a lens for TRY. This is a nono.

      This beam path will be relayed again as the TRY, green beam and IP-ang get there place.

 Oplev is disabled. I removed one of the steering mirrors because it was on the green beam  path.

  8659   Thu May 30 18:29:21 2013 ManasaUpdateGeneralALS medm screens edited

Renaming of the c1gcv model earlier (elog 7011) had left white boxes in most of the ALS medm screens.

Channels names were corrected. No more white boxes.

  8658   Thu May 30 17:18:58 2013 JenneUpdateASCNew POP path

I have placed the G&H mirrors and the Y1 as pictured in my proposed layout in elog 8649.  The distance between the 2" lens and the PDs has increased, so the focus point is all wrong.  I have measured the distances between optics on the table, and will pick new lenses and finish the POP layout later today or tomorrow.

For now, here are the powers measured using the Ophir power meter:

----------------------

PRM-ITMY lock, POPDC was ~190 counts

5.29 uW after Y1 weird angle.  Can't see beam before then to measure

5.27 uW before BS50
3.5 uW before razor PD
3.00 uW before 110PD

 ------------------------

After installing G&H mirrors, replacing BS-98 with Y1:

4.94 uW before y1 (after G&H's)
4.92 uW after y1
2.66 uW before razor PD
1.61 uW before 110PD

  8657   Thu May 30 11:33:26 2013 JamieConfigurationComputer Scripts / ProgramsASS medm/model changes need to be committed to SVN

There are a lot of changes to the ASS stuff that have not been committed to the SVN:

controls@rossa:/opt/rtcds/userapps/release/isc/c1 0$ svn status | grep -v '?'
M       medm/c1als/C1ALS_X_SLOW.adl
D       medm/c1ass/C1ASS_TRY_YAW_LOCKIN.adl
D       medm/c1ass/ASS_SERVOS.adl
D       medm/c1ass/ctrl_yaw_mtrx.adl
D       medm/c1ass/C1ASS_QPDS.adl
D       medm/c1ass/C1ASS_SEN_YAW_MTRX.adl
M       medm/c1ass/C1ASS_XARM_SEN_MTRX.adl
D       medm/c1ass/SITEMODEL_LOCKINNAME.adl
D       medm/c1ass/C1ASS_TRX_YAW_LOCKIN.adl
D       medm/c1ass/C1ASS_LOCKIN1.adl
D       medm/c1ass/C1ASS_LOCKIN2.adl
D       medm/c1ass/C1ASS_LOCKIN3.adl
D       medm/c1ass/C1ASS_LOCKIN4.adl
D       medm/c1ass/C1ASS_LOCKIN5.adl
D       medm/c1ass/C1ASS_LOCKIN6.adl
D       medm/c1ass/C1ASS_LOCKIN7.adl
D       medm/c1ass/C1ASS_LOCKIN8.adl
D       medm/c1ass/C1ASS_LOCKIN9.adl
D       medm/c1ass/C1ASS_REFL11I_PIT_LOCKIN.adl
M       medm/c1ass/C1ASS.adl
D       medm/c1ass/C1ASS_LOCKIN10.adl
D       medm/c1ass/C1ASS_LOCKIN11.adl
D       medm/c1ass/C1ASS_LOCKIN12.adl
D       medm/c1ass/C1ASS_LOCKIN13.adl
D       medm/c1ass/C1ASS_LOCKIN14.adl
D       medm/c1ass/C1ASS_LOCKIN15.adl
D       medm/c1ass/sen_yaw_mtrx.adl
D       medm/c1ass/C1ASS_LOCKIN16.adl
D       medm/c1ass/C1ASS_LOCKIN17.adl
D       medm/c1ass/C1ASS_DOF_YAW.adl
D       medm/c1ass/C1ASS_LOCKIN18.adl
D       medm/c1ass/C1ASS_LOCKIN19.adl
D       medm/c1ass/C1ASS_TRY_PIT_LOCKIN.adl
D       medm/c1ass/ctrl_pit_mtrx.adl
D       medm/c1ass/C1ASS_SEN_PIT_MTRX.adl
D       medm/c1ass/C1ASS_LOCKIN20.adl
D       medm/c1ass/C1ASS_LOCKIN21.adl
D       medm/c1ass/C1ASS_LOCKIN22.adl
D       medm/c1ass/C1ASS_LOCKIN23.adl
D       medm/c1ass/C1ASS_LOCKIN24.adl
D       medm/c1ass/C1ASS_LOCKIN25.adl
D       medm/c1ass/C1ASS_LOCKIN26.adl
D       medm/c1ass/C1ASS_LOCKIN27.adl
D       medm/c1ass/C1ASS_TRX_PIT_LOCKIN.adl
D       medm/c1ass/C1ASS_LOCKIN28.adl
D       medm/c1ass/C1ASS_LOCKIN29.adl
D       medm/c1ass/C1ASS_XARM_QPDS.adl
D       medm/c1ass/C1ASS_YARM_QPDS.adl
M       medm/c1ass/C1ASS_XARM_OUT_MTRX.adl
D       medm/c1ass/ASS_SEN_MTRX.adl
D       medm/c1ass/ASS_LOCKINS.adl
D       medm/c1ass/sen_pit_mtrx.adl
D       medm/c1ass/C1ASS_REFL11I_YAW_LOCKIN.adl
D       medm/c1ass/C1ASS_LOCKIN30.adl
D       medm/c1ass/C1ASS_DOF_PIT.adl
M       models/c1ass.mdl
controls@rossa:/opt/rtcds/userapps/release/isc/c1 0$
  8656   Thu May 30 11:28:34 2013 JamieConfigurationCDSc1als model cleanup

The c1als model was pulling out some ADC0 connections that were no longer used for anything:

  • ADC_0_1 --> sfm "FD" --> IPC "C1:ALS-SCX_FD"
  • ADC_0_5 --> sfm "OCX" --> term
  • ADC_0_6 --> sfm "ADC" --> term

The channels would have shown up as C1:ALS-FD, C1:ALS-OCX, C1:ALS-ADC.  The IPC connection that presumably was meant to go to c1scx is not connected on the other end.

I removed all this stuff from the model and rebuilt/restarted.

  8655   Thu May 30 11:10:00 2013 SteveUpdateendtable upgradeETMY - oplev

Quote:

 Temporary oplev in place. The spot on the qpd is still big. My two lens solution did not work.

I will finalize optical component position of the oplev after the the arm transmitted and green beam optics in place. They have priority.

 

 Oplev spot size on qpd ~ 1 mm 

PS: I realized it later that the returning beam is going through a lens for TRY. This is a nono.

      This beam path will be relayed again as the TRY, green beam and IP-ang get there place.

Attachment 1: ETMYoplev.jpg
ETMYoplev.jpg
  8654   Thu May 30 10:40:59 2013 JamieConfigurationCDSAttempt to cleanup c1ioo ADC connections

I have attempted to reconcile all of the ADC connections to c1ioo.  Upon close inspection, it appears that there was a lot of legacy stuff hanging around.  Either that or things have not been properly connected.

The c1ioo front end machine has two ADC cards, ADC0 and ADC1, which are used by two models, c1ioo and c1als.  The CURRENT ADC connections are listed in the table below.  The yellow cells indicate connections that were moved.  The red cells indicate connections that were removed/unplugged:

  channel block connection channel usage  model
ADC0 8-15 MC WFS1 interface   MC WFS1 c1ioo
16-23 MC WFS2 interface   MC WFS2 c1ioo
0-7

generic interface card (2 pin lemo)

0    
1    
2    
3 ALS TRX c1als
4 ALS TRY c1als
5    
6 MCL c1ioo
7 MCF c1ioo

 

  channel block connection channel usage model
ADC1 0-31 1U interface board 0/1 (J1A) PSL FSS MIXER/NPRO c1ioo
2/3 (J2) ALS BEAT X/Y DC c1als
4/5 (J3) PSL eurocrate DAQ interface J4  
6/7 PSL eurocrate DAQ interface J5  
8/9 PSL eurocrate DAQ interface J6  
10/11 MC eurocrate DAQ interface J1  
12/13 MC servo board DAQ  
14/15 (J8)    
16/17 (J9A) UNLABELLED ("DAQ ISS1"???)  
18/19 (J10) "DAQ ISS2"  
20/21 "DAQ ISS3"  
22/23 ALS BEAT X I/Q c1als
24/25 ALS BEAT Y I/Q c1als
26/27    
28/29    
30/31 (J16)    

The following changes were made:

  • "MC L" had been connected to ADC_0_0, moved to ADC_0_6
  • "MC F" had been connected to ADC_0_6, moved to ADC_0_7

The c1ioo model was rebuilt/restarted to reflect this change.

The PSL-FSS_MIXER and PSL-FSS_NPRO connections were broken in the c1ioo so I fixed them when I moved the MC channels.

All the removed connections from ADC1 were not used by any of the front end models, which is why I unplugged them.  Except for the MC DAQ interface J1 and MC servo DAQ connections, I left all other cables plugged in to wherever they were coming from.  The MC cables I did fully remove.

I don't know what these connections were meant for.  Presumably they expose they expose some useful DAQ channels that we're now getting elsewhere, but I'm not sure.  We don't currently have an ISS, which is presumably why the cables labelled "ISS" are not going anywhere.

TODO

I would like to see some more 4-pin lemo --> double BNC cables made.  That would allow us to more easily use the ADC1 generic interface board:

  • Moved ALS TRX/Y to ADC1, so that we can keep all the ALS connections together in ADC1.
  • POP QPD X/Y/SUM

We should also figure out if we're sub-optimally using the various "DAQ" connections to the DAQ cable connectiosn to the eurocrate DAQ interface cards and servo boards.

  8653   Thu May 30 01:02:41 2013 ManasaUpdateGreen LockingFound it!

X green beat note found!

Key points
1. Near-field and far-field alignment on the PSL table. The near-field alignment checked by looking at the camera and the far-field alignment checked by allowing the beams to propagate by removing the DC PD.
2. Check laser temperature and get a sense of how the offset translates to the actual laser temperature.
3. Get an idea of the expected temperature of laser using the plot in elog.

Data
PSL laser temperature = 31.45 deg C

X end laser temperature = 39.24 deg C
C1-ALS-X_SLOW_SEERVO2_OFFSET = 4810
Amplitude of beat note = -40dBm

I do not understand why
1. The amplitude of beatnote falls linearly with frequency (peak traced using 'hold' option of the spectrum analyzer).
2. I found the beat note at the RF output of the PD. Earlier, while I was trying to search for the beatnote from the RFmon output of the betabox, there was a strong peak at 29.6MHz that existed even when the green shutters were closed. It's source has to be traced.

Next
Solve beatbox puzzle and lock arm using ALS.

IMG_0598.JPGIMG_0599.JPG

  8652   Tue May 28 22:11:23 2013 AnnalisaUpdate40m upgradingETMY - Yarm shutter cable

 

 For some strange reason the Yarm shutter cable runs up to the POY table, where it is connected to another cable going to the rack. It has to be put off from the table, at least. It would be better to have only one cable going directly to the rack.

I roughly aligned the green into the Yarm and I've seen the green beam flashing on the PSL table, but the mode matching is not so good and I get an higher order mode, so I'm going to fix the mode matching tomorrow.

ELOG V3.1.3-